OpenStudy (anonymous):

If f(x) = 8 – 3|x + 2|, find f(–6) A. –4 B. 20 C. 4 D. –16

OpenStudy (zzr0ck3r):

f(x) = 8 – 3|x + 2| step at a time we will plug in -6 for x and get -6+2 = -4 now we do |-4| = 4 now we do 3*4 = 12 now we do 8-12 = -4 in all \(f(-6) = 8-3|-6+2|=8-3|-4| = 8-3*4 = 8-12 = -4\)

OpenStudy (zzr0ck3r):

you should be able to do the next one, do it step at a time, and start inside the ||